a pole is broken at a height of 6m from the ground and its top touches the ground at a distance of 8m from the base of the tree find the original height of the tree
Answer by fractalier(6550)

If you draw the diagram you have a right triangle with 6m going up and 8m at the base.
Using the Pythagorean Theorem, we can find the length of the fallen part of the pole...
c^2 = a^2 + b^2
c^2 = 6^2 + 8^2
c^2 = 36 + 64
c^2 = 100
c = 10
Thus when the pole is fixed and upright again, its total length is 16 m.